# Constant-Memory Strategies in Stochastic Games: Best Responses and Equilibria

### v1

@@ 4.1 Mixed Strategies Are Something More Than Behavioral Strategies
$\begin{split}V_{mix}(S)&=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\sum_{\iota}p_{\iota}\cdot Q_{\mathcal{M}(\pi^{\iota}_{-i})}(S,a_{i})\\ &=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\sum_{\iota}p_{\iota}\cdot\Big{[}R_{\pi^{\iota}_{-i}}(S,a_{i})+\gamma\sum_{S^{\prime}}T_{\pi^{\iota}_{-i}}(S^{\prime}|S,a_{i})V_{\mathcal{M}(\pi^{\iota}_{-i})}(S^{\prime})\Big{]}\\ &=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\sum_{\iota}p_{\iota}\cdot\Big{[}\sum_{a_{i}\in\mathcal{A}_{-i}}R_{i}(S,a)\pi^{\iota}_{-i}(a_{-i}|S)+\gamma\sum_{S^{\prime}}\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\pi^{\iota}_{-i}(a_{-i}|S)V_{\mathcal{M}(\pi^{\iota}_{-i})}(S^{\prime})\Big{]}\\ &=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\Big{[}\sum_{a_{i}\in\mathcal{A}_{-i}}R_{i}(S,a)\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)+\gamma\sum_{S^{\prime}}\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)V_{\mathcal{M}(\pi^{\iota}_{-i})}(S^{\prime})\Big{]}\\ &{\color[rgb]{0,.5,.5}\ \neq\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\Big{[}\sum_{a_{i}\in\mathcal{A}_{-i}}R_{i}(S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\vec{p})}}+\gamma\sum_{S^{\prime}}\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\vec{p})}}V_{mix}(S^{\prime})\Big{]}}\\ \end{split}$
@@
$\begin{split}V_{beh}(S)=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\Big{[}&\sum_{a_{i}\in\mathcal{A}_{-i}}R_{i}(S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\vec{p})}}\\ +&\gamma\sum_{S^{\prime}}\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\vec{p})}}V_{beh}(S^{\prime})\Big{]}\end{split}$ (13)
@@
Remark 1 (Singleton state but positive-length memory).
As both increasing the length of memory and the number of environment states will lead to a multi-state MDP from each individual agent’s perspective, a natural corollary is that utility equivalence between a mixed strategy and its induced behavioral strategy does not necessarily hold for $K$-memory strategies once $K$ is positive, even under repeated games.
One may further wonder whether a group of agents can form some equilibrium if all of them play mixed strategies, i.e., $\{(\Pi^{K+}_{i},\vec{p}_{i})\}_{i\in\mathcal{N}}$. The story is that, if the support $\{\Pi^{K+}_{i}\}_{i\in\mathcal{N}}$ and a distribution over initial states $d_{0}\in\Delta(\mathcal{S})$ can be specified in the first place, the stochastic game can be further reduced to a normal-form game $\langle\mathcal{N},\{\Pi^{K+}_{i}\}_{i\in\mathcal{N}},\{u_{i}\}_{i\in\mathcal{N}}\rangle$ (with $u_{i}$ slighted overloaded) as follows,
  1. 1.
@@
Under this sense and provided that the reduced game is finite, invoking Nash’s well-known existence theorem, we can conclude that there must exists a mixed strategy NE $\{\vec{p_{i}}^{*}\}_{i\in\mathcal{N}}$. That is, given fixed supports $\{\Pi^{K+}_{i}\}_{i\in\mathcal{N}}$, no one will be strictly better off by unitarily deviating from $\{\vec{p_{i}}^{*}\}_{i\in\mathcal{N}}$ to another distribution for mixing over its support strategies. However, the application of this result remains an open (and perhaps even unjustified) problem. One idea might be promising: as we will see later, finding a behavioral strategy best response to a mixed strategy is computationally hard, but will it helps if it allows for finding a mixed strategy best response instead?
4.2 Computing Best Responses to Mixed Strategies is Intractable
We will first show that computing the best response against a mixed $K$-memory strategy reduces to optimally solving an infinite-horizon partially observable MDPs (POMDPs), so that the complexity/computability results [31, 32] can naturally follow. It turns out the reduced ones belong to a subclass of generic POMDPs, namely Contextual MDP (CMDPs), although it may not potentially imply less challenging computation. To show that such reduction does not complicate the original problem, we also construct a reduction from the computational problem of solving CMDPs back to that of computing best responses (against mixed strategies).
@@ 4.2 Computing Best Responses to Mixed Strategies is Intractable
Given a mixed strategy profile $(\Pi^{K+}_{-i},\vec{p})$ of the opponents, computing the best response for agent $i$ can be reduced to optimally solving an infinite-horizon POMDP.
@@
We first construct the reduction to the corresponding POMDP. The POMDP is given as a tuple $\langle\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i},\mathcal{A}_{i},\mathcal{H}^{K}\times\mathcal{S},\mathbf{T},\mathbf{O},\mathbf{R},\gamma\rangle$,
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    $\mathbf{T}:(\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i})\times\mathcal{A}_{i}\mapsto\Delta(\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i})$ denote transition function, mathematically defined as
  5. 5.
    $\mathbf{O}\Big{(}(H,S,\pi_{-i}^{\prime})\Big{)}=(H,S)$
  6. 6.
    $\mathbf{R}:(\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i})\times\mathcal{A}_{i}\mapsto\mathbbm{R}$, mathematically defined as
  1. 1.
      $\{\vec{q},S_{0},a_{i,0},a_{-i,0},S_{1},a_{i,1},a_{-i,1},\cdots,S_{t}\}$
      and that in the reduced POMDP is all the historic observations
      $\begin{split}&\{\vec{q},S_{0}\}\\ \cup&\{(S_{0},a_{i,0},a_{-i,0}),S_{1}\}\\ \cup&\{(S_{0},a_{i,0},a_{-i,0}),(S_{1},a_{i,1},a_{-i,1}),S_{2}\}\\ &\cdots\\ \cup&\{(S_{t-K},a_{i,t-K},a_{-i,t-K}),\cdots,(S_{t-1},a_{i,t-1},a_{-i,t-1}),S_{t}\}\\ =&\{\vec{q},S_{0},a_{i,0},a_{-i,0},S_{1},a_{i,1},a_{-i,1},\cdots,S_{t}\}\end{split}$
  2. 2.
  3. 3.
One can see that the reduced POMDPs in the above theorem actually belong to a subclass of generic POMDPs, where a state consists of variables that can be directly observed and other hidden ones. Such a subclass is specially termed as Mixed observability MDPs (MOMPDs) [36, 6, 28]. It has been shown that planning algorithms originally developed for POMDPs are significantly faster for those factorized models like MOMDPs in practice. A more restricted model is the Contextual MDP (CMDP) [23, 11], which can be viewed as a special case of MOMDP where there is no transition between different hidden state variables. Although one can see that CMDP $\subseteq$ MOMDP $\subseteq$ POMDP in terms of computational hardness, the complexity results in general for the former two remain open. So far, the common conjecture is that neither CMDP nor MOMDP is significantly easier to solve than POMDP, and it is proven that optimally solving infinite-horizon POMDPs is undecidable [32].
This result highly pertains to some discussions on type-based methods for single-agent planning under the presence of multiple other agents [3, 4, 2, 55]. Albrecht and Ramamoorthy [3] characterized the general problem from a conceptual perspective, where each opponent’s strategy acts as an oracle that can be queried, but left the concrete issues regarding implementation unresolved. As a supplementary, Zhu and Lin [55] provided a spectrum of implementable planners for the stationary base case, i.e., each support strategy of the opponent’s mixed strategy is a stationary one. Here, Theorem 7 further generalizes it to constant-memory strategies, allowing all the formulations in [55] to be carried over to the entire family of constant-memory strategies..

### v2

@@ A.3. For Theorem 10
[…]
@@
$\begin{split}V_{beh}(S)=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\Big[\sum_{a_{i}\in\mathcal{A}_{-i}}R_{i}(S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\@vec{p})}}+\gamma\sum_{S^{\prime}}\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\@vec{p})}}V_{beh}(S^{\prime})\Big]\end{split}$ (17)
@@
A.4. For Theorem 11
Proof.
Recall the corresponding POMDP is given as the tuple $\langle\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i},\mathcal{A}_{i},\mathcal{H}^{K}\times\mathcal{S},\mathbf{T},\mathbf{O},\mathbf{R},\gamma\rangle$,
  1. (1)
  2. (2)
  3. (3)
  4. (4)
[…]
  5. (5)
    $\mathbf{O}\Big((H,S,\pi_{-i}^{\prime})\Big)=(H,S)$
  6. (6)
    $\mathbf{R}:(\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i})\times\mathcal{A}_{i}\mapsto\mathbb{R}$, mathematically defined as
  1. (1)
      $\{\@vec{q},S_{0},a_{i,0},a_{-i,0},S_{1},a_{i,1},a_{-i,1},\cdots,S_{t}\}$
      and that in the reduced POMDP is all the historic observations
      $\begin{split}&\{\@vec{q},S_{0}\}\\ \cup&\{(S_{0},a_{i,0},a_{-i,0}),S_{1}\}\\ \cup&\{(S_{0},a_{i,0},a_{-i,0}),(S_{1},a_{i,1},a_{-i,1}),S_{2}\}\\ &\cdots\\ \cup&\{(S_{t-K},a_{i,t-K},a_{-i,t-K}),\cdots,(S_{t-1},a_{i,t-1},a_{-i,t-1}),S_{t}\}\\ =&\{\@vec{q},S_{0},a_{i,0},a_{-i,0},S_{1},a_{i,1},a_{-i,1},\cdots,S_{t}\}\end{split}$
  2. (2)
  3. (3)
Appendix B Nash Equilibria for Mixed Constant-memory Strategies
We examine whether a group of agents can form some equilibrium if all of them play mixed strategies, i.e., $\{(\Pi^{K+}_{i},\@vec{p}_{i})\}_{i\in\mathcal{N}}$. The story is that, if the support $\{\Pi^{K+}_{i}\}_{i\in\mathcal{N}}$ and a distribution over initial states $d_{0}\in\Delta(\mathcal{S})$ can be specified in the first place, the stochastic game can be further reduced to a normal-form game $\langle\mathcal{N},\{\Pi^{K+}_{i}\}_{i\in\mathcal{N}},\{u_{i}\}_{i\in\mathcal{N}}\rangle$,
  1. (1)
@@ Appendix B Nash Equilibria for Mixed Constant-memory Strategies
Under this sense and provided that the reduced game is finite, invoking Nash’s well-known existence theorem, we can conclude that there must exists a mixed strategy NE $\{\@vec{p_{i}}^{*}\}_{i\in\mathcal{N}}$. That is, given fixed supports $\{\Pi^{K+}_{i}\}_{i\in\mathcal{N}}$, no one will be strictly better off by unitarily deviating from $\{\@vec{p_{i}}^{*}\}_{i\in\mathcal{N}}$ to another distribution for mixing over its support strategies. However, the application of this result remains an open (and perhaps even unjustified) problem. One idea might be promising: as we will see later, finding a behavioral strategy best response to a mixed strategy is computationally hard, but will it helps if it allows for finding a mixed strategy best response instead?
Appendix C Details for the Iterated Prisoner’s Dilemma
For the readers’ convenience, we first echo the payoff matrix of the Prisoner’s Dilemma in Table 1, with $T>R>P>S$.
C D
C (R, R) (S, T)
D (T, S) (P, P)
Table 1. The payoff matrix of the Prisoner’s Dilemma
C.1. Some Formal Results
@@
Given the opponent playing a “NN-Tit(s)-for-$M$-Tat(s)” strategy, there exists a best response strategy that can be implemented with (M1)(M-1)-memory.
@@
As a “NN-Tit(s)-for-$M$-Tat(s)” strategy is a $\max(N,M)$-memory strategy, then Theorem 2 implies that there must exist a $\max(N,M)$-memory strategy serving as a best response. We then show that 1) there exists a strategy within $\max(N,M)$-memory that can result in the following payoff sequence; and 2) any strategy that can result in the following payoff sequence is a best response,
$\underbrace{(T,S),(T,S),\cdots,(T,S)}_{M-1},(R,R),\underbrace{(T,S),(T,S),\cdots,(T,S)}_{M-1},(R,R),\cdots$
One can construct a $M$-memory strategy as “$X$-D(s)-before-one-C”, with X=M1X=M-1 here. By its name, it means to start with defection, and then cooperate only after (M1)(M-1) defections. As the opponent will retaliate only when being defected $M$ times, therefore, the constructed strategy will make the opponent cooperate all the time, hence the above payoff sequence. In fact, implementing such a “(M1)(M-1)-D(s)-before-one-C” strategy only requires the agent to keep track of the past (M1)(M-1) actions of its own: if it has played one defection in the past (M1)(M-1) rounds, then keep cooperating, otherwise defect for one round. Hence, it is actually a (M1)(M-1)-memory strategy.
Also, to have a better sequence, one should note that, it is impossible to “flip” every $(R,R)$ to $(T,S)$, as the opponent will definitely defect after the $M$-the defection. The only way to better off is to flip some of the $(R,R)$’s to $(T,S)$’s without sacrificing too much of the future return. In fact, it is only possible to flip the first $(R,R)$ to $(T,S)$ (i.e., by playing an “All-D” strategy), and the rest all will be changed to $(P,P)$. As we will show detailed calculations in the next subsection, when the agent is patient enough (i.e., with high discount factor), such a deviation is not profitable. Nevertheless, one should note that even when the agent is impatient, and therefore, adopts the “All-D” strategy, the strategy can be implemented with 0-memory. ∎
C.2. Phase Transition
We also mentioned that the best response will transit from a “$X$-D(s)-before-one-C” strategy to a “All-D” one, when the discounted factor keeps decreasing (i.e., the agents become less patient and more myopic). Now we formally derive the critical point of the discount factor that triggers such phase transition.
Assume the column player is playing a “NN-Tit(s)-for-$M$-Tat(s)” strategy, we compute the discounted accumulated payoffs of the row player performing different responding strategies.
  1. (1)
    When the row player plays a “(M1)(M-1)-D(s)-before-one-C” strategy, the payoff sequence will be
    $\underbrace{(T,S),(T,S),\cdots,(T,S)}_{M-1},(R,R),\underbrace{(T,S),(T,S),\cdots,(T,S)}_{M-1},(R,R),\cdots$
    The discounted accumulated payoff will be
    $\begin{split}&\big(T+\gamma T+\cdots+\gamma^{M-2}T+\gamma^{M-1}R\big)+\big(\gamma^{M}T+\gamma^{M+1}T+\cdots+\gamma^{2M-2}T+\gamma^{2M-1}R\big)+\cdots\\ =&\frac{T}{1-\gamma^{M}}+\frac{\gamma T}{1-\gamma^{M}}+\cdots+\frac{\gamma^{M-2}T}{1-\gamma^{M}}+\frac{\gamma^{M-1}R}{1-\gamma^{M}}\\ =&\frac{\frac{T}{1-\gamma^{M}}(1-\gamma^{M-1})}{1-\gamma}\end{split}$ (18)
  2. (2)
    When the row play plays an “All-D” strategy, the payoff sequence will be
    $\underbrace{(T,S),(T,S),\cdots,(T,S)}_{M},\underbrace{(P,P),\cdots}_{\text{forever}}$
    The discounted accumulated payoff will be
    $\begin{split}&\big(T+\gamma T+\cdots+\gamma^{M-2}T+\gamma^{M-1}T\big)+\big(\gamma^{M}P+\gamma^{M+1}P+\cdots\big)\\ =&\frac{T(1-\gamma^{M})}{1-\gamma}+\frac{\gamma^{M}P}{1-\gamma}\end{split}$ (19)
Compare Eq (18) and Eq (19), we first simplify it to the following, and then solve $\gamma$ in terms of the other constants.
$T(1-\gamma^{M})(1-\gamma^{M})+P\gamma^{M}(1-\gamma^{M})=T(1-\gamma^{M-1})+R\gamma^{M-1}(1-\gamma)$ (20)
It is intractable to solve it manually. In fact, such an equation with its order being a variable is infeasible to solve even resorting to sophisticated libraries like SymPy. [^4] https://www.sympy.org/en/index.html. Therefore, we substitute $M$ with concrete values first and then solve $\gamma$ using SymPy.
We list the closed-form solutions for $M$ up to 3, and substitute {T=2, R=1, P=0, S=-1} to the final expression.
  1. (1)
    When $M=1$.
    $\begin{split}\gamma&=\frac{R-T}{P-T}=\frac{1}{2}\end{split}$
    The other solution is 1.
  2. (2)
    When $M=2$.
    $\gamma=\frac{-P+T-\sqrt{(P^{2}+4PR-6PT-4RT+5T^{2})}}{2P-2T}=-\frac{1}{2}+\frac{\sqrt{3}}{2}\approx 0.366025$
    The other three solutions are 0, 1, and a negative real number.
  3. (3)
    When $M=3$.
    $\begin{split}\gamma&=\frac{-\Big(\frac{1}{2}\sqrt{(-7+\frac{27(-R+T)}{(P-T)})^{2}+32}-\frac{7}{2}+\frac{27(-R+T)}{2(P-T)}\Big)^{\frac{1}{3}}}{3}-\frac{1}{3}+\frac{2}{3\Big(\frac{1}{2}\sqrt{(-7+\frac{27(-R+T)}{(P-T)})^{2}+32}-\frac{7}{2}+\frac{27(-R+T)}{2(P-T)}\Big)^{\frac{1}{3}}}\\ &=-1/3-(-41/4+3*\sqrt{201}/4)^{1/3}/3+2/(3*(-41/4+3*\sqrt{201}/4)^{1/3})\\ &\approx 0.342508\end{split}$
    The other four solutions are 0, 1, and two complex numbers.
C.3. Sample Outputs of the Computed Best Responses
We here present the computed best response for Player 1 against Player 2 who plays a “2-Tits-For-2-Tats” strategy.
BR to [2 Tits For 2 Tats], val = 15.26 |
| Histories    | P1 action   | P2 action   |   P1 val |
| []           | D           | C           |  15.2632 |
| [D, D]       | C           | C           |  14.7368 |
| [D, C]       | C           | C           |  14.7368 |
| [C, D]       | D           | C           |  15.2632 |
| [C, C]       | D           | C           |  15.2632 |
| [D, D, D, D] | C           | D           |  11.7368 |
| [D, D, D, C] | C           | D           |  11.7368 |
| [D, D, C, D] | D           | C           |  15.2632 |
| [D, D, C, C] | D           | C           |  15.2632 |
| [D, C, D, D] | C           | D           |  11.7368 |
| [D, C, D, C] | C           | D           |  11.7368 |
| [D, C, C, D] | D           | C           |  15.2632 |
| [D, C, C, C] | D           | C           |  15.2632 |
| [C, D, D, D] | C           | C           |  14.7368 |
| [C, D, D, C] | C           | C           |  14.7368 |
| [C, D, C, D] | D           | C           |  15.2632 |
| [C, D, C, C] | D           | C           |  15.2632 |
| [C, C, D, D] | C           | C           |  14.7368 |
| [C, C, D, C] | C           | C           |  14.7368 |
| [C, C, C, D] | D           | C           |  15.2632 |
| [C, C, C, C] | D           | C           |  15.2632 |
Appendix D Details for the Iterated Traveler’s Dilemma
D.1. Generalized (One-Shot) Traveler’s Dilemma
It is conventionally a two-player game, but here we introduce a generalized multi-player version and then present the two-player version as a special case. This domain is of great significance as one can see its connection with PD, auction with the same common value and negotiation. A one-shot multiple-player Traveller’s Dilemma (TD) consists of three parameters, denoted as $TD(N,k,A)$, where $N=[1..n]$ is the set of $n$ participating agents, $k>1$ is a constant coefficient, and $A\subseteq\mathbb{N}$ is a finite set of possible (non-negative) biddings. Given a bidding profile $\@vec{a}=(a_{i},a_{-i})\in A^{N}$ that is simultaneously reported from all agents, the utility for agent $i$ is calculated as
$u_{i}(a_{i},a_{-i})=\min(a_{i},a_{-i})+k\cdot sign(\min(a_{-i})-a_{i})$
where we slightly abuse $\min$ by allowing it to first flatten all its arguments which might be a scaler or a vector and then return whichever element that is the minimum.
Figure 4. The payoff matrix for $TD(2,2,[2..100])$. The figure is borrowed from its wiki page.
Profitable deviation.
Suppose $a_{i}=\alpha$ and $a_{-i}=\alpha\cdot\mathbf{1}$, then $u_{i}(a_{i},a_{-i})=u_{i}(\alpha,\alpha\cdot\mathbf{1})=\alpha$. Consider possible deviation as $a_{i}^{\prime}\leftarrow\alpha-x$, the utility will be changed to $u_{i}(a_{i}^{\prime},a_{-i})=\alpha-x+k$. As long as there exists an $x$ such that $\alpha-x\in A$ and $k>x$, this will be a profitable deviation. Note that one would never deviate to higher bids, say $a_{i}^{\prime\prime}\leftarrow\alpha+y\in A$, $u_{i}(a_{i}^{\prime\prime},a_{-i})=\alpha-k<\alpha=u_{i}(a_{i},a_{-i})$. In a nutshell, everyone is incentivized to bid slightly lower than the current lowest one.
Nash equilibrium.
There is a unique NE, $a_{1}=\cdots=a_{n}=\min(A)$, as none can bid even lower. Note that in terms of better response instead of best response, one does not have to bid lower than the current lowest one, she can just bid the same as the current lowest one. For example, $u_{i}(\alpha,(\alpha-1)\cdot\mathbf{1})=\alpha-1-k$, and $u_{i}(\alpha-1,(\alpha-1)\cdot\mathbf{1})=\alpha-1$.
Two-player traveller’s dilemma.
Figure 4 is an instance of $TD(N=2,k=2,A=[2..100])$. For a detailed story for this game, please refer to the wiki page [^5] https://en.wikipedia.org/wiki/Traveler%27s_dilemma. In this $TD(2,2,[2..100])$ game, each agent required to bid an integer in the interval of $[2..100]$. The utility can be rewritten as
$\begin{cases}u_{1}(a_{1},a_{2}):=\min(a_{1},a_{2})+2\cdot sign(a_{2}-a_{1})\\ u_{2}(a_{1},a_{2}):=\min(a_{1},a_{2})+2\cdot sign(a_{1}-a_{2})\end{cases}$
Potentials.
We show that TD has a generalized ordinal potential by introducing the following definitions first.
Definition 0 (Generalized ordinal potential).
A function $P:A^{N}\mapsto\mathbb{R}$ is called a generalized ordinal potential, if
$\forall i\in N,\forall a_{-i}\in A^{N-1},\forall a_{i},a_{i}^{\prime}\in A,u_{i}(a_{i},a_{-i})-u_{i}(a_{i}^{\prime},a_{-i})>0\implies P(a_{i},a_{-i})-P(a_{i}^{\prime},a_{-i})>0$
We first show the above $TD(2,2,[2..100])$ has a generalized ordinal potential as a warming up example, and then extend it to the general case.